### v2

# JingFang: An Expert-Level Large Language Model for Traditional Chinese Medicine Clinical Consultation and Syndrome Differentiation-Based Treatment

Abstract
The effective application of traditional Chinese medicine (TCM) requires extensive knowledge of TCM and clinical experience. The emergence of Large Language Models (LLMs) provides a solution to this, while existing LLMs for TCM exhibit critical limitations of incomplete clinical consultation and diagnoses, as well as inaccurate syndrome differentiation. To address these issues, we establish JingFang (JF), a novel TCM LLM that demonstrates the level of expertise in clinical consultation and syndrome differentiation. We propose a Multi-Agent Collaborative Chain-of-Thought Mechanism (MACCTM) for comprehensive and targeted clinical consultation, enabling JF with effective and accurate diagnostic ability. In addition, a Syndrome Agent and a Dual-Stage Recovery Scheme (DSRS) are developed to accurately enhance the differentiation of the syndrome and the subsequent corresponding treatment. JingFang not only facilitates the application of LLMs but also promotes the effective application of TCM for healthcare.
1 Introduction
In recent years, the emergence of Large Language Models(LLMs) such as GPT-4o[1, 18], Llama3 [12], Deepseek-MOE [7],Deepseek v3 [10], Gemini-1.5-pro [23] and Qwen2.5 [30] has significantly advanced the field of natural language processing. These models not only possess exceptional language comprehension and generation capacity [33], but also demonstrate strong cross-domain adaptability [19, 3, 25] by integrating techniques such as pre-training [15], fine-tuning [11], reinforcement learning [4], human feedback alignment [32], and knowledge graph retrieval [27]. Among the numerous cross-domain applications of LLM, TCM has become one of the key focuses due to its importance for health care and disease treatment [20, 34, 21].
There are several studies have made some meaningful attempts, e.g., HuaTuo [26](a TCM diagnostic system based on LLM that focuses on simulating the syndrome differentiation thinking process of TCM physicians), TCMChat [8](the generative LLM for TCM built on Baichuan2-7B-Chat and trained on 6 task categorie diverse TCM data), Zhongjing [31] (the reinforcement learning with human feedback combines with CMtMedQA boosted the model’s proactive reasoning and knowledge alignment in complex TCM diagnostics), HuaTuo [5](It proposes an innovative one-stage domain adaptation protocol that unifies heterogeneous data into instruction-output pair format and introduces a data priority sampling strategy), ShenNong [28](It proposes a framework that improves large language models’ performance on TCM tasks using limited data through a two-stage approach: supervised fine-tuning with medical cases followed by reinforcement learning from AI feedback), and SunSimiao [29](the 7B model is fine-tuned from Qwen2-7B with high-quality medical data, achieving excellent performance equivalent to 30B-scale models on CMB-Exam).
However, the critical limitations of the effective application of LLM in TCM still remain. The core characteristic of TCM is the emphasis on syndrome differentiation and treatment, which requires physicians to conduct a comprehensive clinical consultation, accurately analyze patients’ symptoms and signs, and then carry out the corresponding treatment. Comprehensive clinical consultation is the foundation for syndrome differentiation and treatment, which largely relies on multiple rounds of professional and targeted inquiries to accurately collect information. The scarcity of large-scale and accurately labeled multi-round consultation data leads to not good enough performance of existing models in clinical consultation and diagnosis. They also lack explicit control over the clinical consultation process and diagnostic reasoning, affecting the interpretability and reliability of the models. In addition, existing models have significant limitations in the precise differentiation of TCM syndromes, which affects their effectiveness in disease treatment and their application in real clinical scenarios.
To address these, the main contributions of this work are summarized as:
  1. 1.
    We develop JF, a TCM LLM that addresses the limitations of existing models and empowers effective applications of LLM in TCM. JF’s innovative framework endows it with expert-level clinical consultation and syndrome differentiation capabilities.
  2. 2.
    The Multi-Agent Collaborative Chain-of-Thought Mechanism (MACCTM) is designed to enhance model’s comprehensive clinical consultation abilities, enabling professional and targeted clinical interactions while strengthening its explicit reasoning, decision and interpretability during the consultation process.
  3. 3.
    The Syndrome Agent has been systematically trained based on preprocessed and structured TCM knowledge, combining with the developed Dual-Stage Retrieval Scheme (DSRS) to substantially improve the model’s capabilities in precise syndrome differentiation and the corresponding treatment.
Figure 1: JF framework, including three main modules: TCM Consultation, TCM Syndrome Differentiation, and TCM Treatment.
2 Method
2.1 Framework of JingFang
As shown in Figure 1, integrating the technology of LLM Agent  [17], Chain-of-Thought (CoT) [6], and Retrieval-Augmented Generation (RAG) [14], the framework of JF consists of three main modules: TCM Consultation, TCM Syndrome Differentiation, and TCM Treatment Reconmmendation, following the authentic process of TCM diagnosis and treatment. The MACCTM in TCM Consultation enables JF’s ability to provide comprehensive and targeted clinical consultation, the Syndrome Agent in TCM Syndrome Differentiation improves JF’s ability to accurately differentiate syndrome, and the DSRS in TCM Treatment Reconmmendation ensures the model to provide accurate treatment recommendations according to diagnosis and syndrome differentiation. See A in details.
2.2 Multi-Agent Collaborative Chain-of-Thought Mechanism
After an in-depth study of the TCM consultation process, we found that experienced TCM experts typically pre-set one or two follow-up questions based on the patient’s known situations, and make adaptive and targeted inquiries to collect more information during the consultation process. Inspired by this, the MACCTM is developed to better fit the authentic TCM consultation process. The prompts: $prompt_{[\ ]}$ here, are particularly designed to facilitate agents to implement a specific function, so the MACCTM can be formulated as follows:
2.2.1 TCM Expert Team Construction
For comprehensive and targeted medical consultation, an Expert Agents Database consisting of TCM General Agents and TCM Specialist Agents from various fields including TCM internal medicine, surgery, gynecology, pediatrics, and so on, is constructed before medical consultation.
[…]
$SD_{i}=f_{LLM}(M,\ K_{ed_{i}},\ prompt_{sd}).$ (1)
[…]
$Team(M)=\{SD_{i},\ GD\},$ (2)
which ensures both the pertinence and comprehensiveness of the consultation. The expert team for consultation is established only once during the consultation process for each patient.
2.2.2 Consultation CoT Construction
To start the CoT construction at the t-$th$ round of consultation, the TCM Medical Record Agent summarizes the patient’s current condition $Con_{t}$ from the current consultation result that is defined as: $C_{t}=\{(Q_{k},A_{k})\mid k\in[0,t]\}$, where $Q_{k}$ and $A_{k}$ is the question and the corresponding answer at k-$th$ round of consultation. Thus, $Con_{t}$ can be defined as:
$Con_{t}=Agen_{rm}(C_{t},\ prompt_{rm}).$ (3)
Each expert in the consultation expert team prioritizes 1-2 subsequent consultation CoT questions with explanations based on the current consultation result $C_{t}$, offering theoretical guidance for future consultations. Specifically, according to $C_{t}$, the TCM Specialized Agent $SD_{i}$ utilizes an LLM in conjunction with professional knowledge and specific prompts $prompt_{qa}$to generate a consultation CoT $qa_{it}$at t-$th$ round of consultation:
$qa_{it}=CoT_{sp}(C_{t},\ SD_{i},\ prompt_{qa})$ (4)
Meanwhile, the TCM General Agent $GD$ incorporates the TCM general knowledge to generate a comprehensive consultation CoT $ga$:
$ga_{it}=CoT_{gen}(C_{t},\ GD,\ prompt_{ga})$ (5)
The expert team’s consultation CoT is constructed by combining $qa_{it}$ and $ga_{it}$ at t-$th$ consultation:
$TCoT_{t}=(qa_{it},\ ga_{it})$ (6)
This consultation CoT lays the foundation for subsequent consultation process. In the following steps, the expert team iteratively revises and optimizes $TCoT_{t}$ to make the consultation process more professional and targeted.
2.2.3 Consultation CoT Integration and Evaluation
Integrating Ten Questions Song [16] (TQS), the widely used classical framework in TCM consultation, we establish a Consultation CoT Evaluation Algorithm (CCEA).
Algorithm 1 Consultation CoT Evaluation Algorithm
1:  Initialization: $scored\_consultation\_cot\leftarrow\emptyset$
2:  for $question$ in ${RCoT_{tf}}$ do
[…]
4:     for item in $ten\_questions\_song$ do
5:        $score\leftarrow sim(emb(question),\ emb(item))$
6:        if $score>com\_score$ then
7:            $com\_score\leftarrow score$
8:        end if
9:     end for
10:     for $item$ in ${core\_questions,medical\_record}$ do
11:        $score\leftarrow sim(emb(question),\ emb(item))$
12:        if $score>per\_score$ then
13:            $per\_score\leftarrow score$
14:        end if
15:     end for
16:     $total\_score\leftarrow com\_score+per\_score$
17:     Store: Add $question$ with its $total\_score$ to $scored\_consultation\_cot$
18:  end for
19:  Return: $scored\_consultation\_cot$
We calculate the similarity of questions in the consultation CoT with those in the TQS as the score of comprehensiveness via an embedding model. For the evaluation the pertinence of the questions, the key issues of each expert (different TCM experts focus on the characteristic diseases and diagnostic points related to their specialty) and the patient’s known condition are incorporated to calculate their similarity as the basis for the pertinence score. The final score is obtained by aggregating the comprehensiveness score and the pertinence score as the evaluation the consultation. Thus, based on the patient’s current condition $Con_{t}$, an Evaluation Agent is designed to conclude the expert team’s consultation CoT $TCoT_{t}$ and evaluate the consultation questions via the TCM knowledge-based CCEA. Thus, a holistic consultation CoT $RCoT_{t0}$ that includes evaluation and summary for each question is generated:
$RCoT_{t0}=CoT_{sum}(TCoT_{t},\ Con_{t},\ prompt_{r})$ (7)
2.2.4 Consultation CoT Analysis and Optimization
Agents in the expert team further analyze the $RCoT_{t0}$ based on the evaluation scores, and provide feedback on whether they agree with it or not. If an expert disagrees, it will provide modification suggestions for low-score questions and optimize the consultation CoT to get a higher score. Through this optimization and feedback process, the TCM expert team ensures that the consultation CoT is both professional and targeted. Specifically, in $j$-th round of feedback, expert suggestions are recorded as $Mod_{j}$ and the $RCoT_{tj}$ is updated as follows:
$RCoT_{tj}=CoT_{opt}(RCoT_{t(j-1)},Mod_{tj},prompt_{r})$ (8)
The modification and optimization process continues until all experts agree on the summary or the maximum number of feedback rounds is reached, a final Consultation CoT $RCoT_{tf}$ is generated to start the formal consultation in this round.
2.2.5 Multi-Round Consultation
Based on the final consultation CoT $RCoT_{tf}$ and the current consultation result $C_{t}$ in the t-$th$ round, a Consultation Agent is developed following the real-world consultation scenarios to generate the next round of consultation question and interact with the patient:
$Q_{t+1}=Q_{gen}(RCoT_{tf},\ Con_{t},\ prompt_{q}).$ (9)
After this, the Medical Record Agent is called to systematically integrate the patient’s condition into medical record $Rec_{t}$, which is
$Rec_{t}=Agen_{in}(M,RCoT_{tf},Con_{t},prompt_{in}).$ (10)
As the process continues, a series of professional and targeted questions is generated to further explore the patient’s condition. When the key information collected in the medical record is completed or the maximum number of consultation rounds is reached according to the Consultation Information Collection Algorithm2, the Medical Record Agent presents the final medical case $Rec_{f}$ to the Syndrome Differentiation Module and Treatment Recommendation Module.
Figure 2: An example of MACCTM in the medical consultation
Fig.2 exhibits an example of how this MACCTM works in a medical consultation scenario. For more detailed and specific examples of the MACCTM in the real medical consultation scenario, see B in the appendix.
2.3 TCM Syndrome Agent
In TCM, the accuracy of syndrome differention is fundamental for precise treatment that requires the high-quality datasets for LLM training. We found that real-world TCM records contain substantial noise resistant to conventional regularization. This noise includes irrelevant administrative statements, such as "patient visited for integrated treatment", and western medical examination references, such as "patient refused cranial CT", which distract models from essential disease information. Additionally, instrument-based diagnostic content, e.g., "blood amylase tests normal, abdominal CT shows colon wall thickening" must be filtered out to maintain focus on TCM-relevant symptom patterns.
In real TCM scenarios, TCM experts usually utilize TQS to comprehensively understand the patient’s symptoms, medical history, and physical signs through a series of fundamental questions, obtaining key information for precise diagnosis. Inspired by this, we propose a universal data preprocessing method by utilizing LLM to automatically extract key information related to the TQS from raw data and strictly retain core information closely related to the patient’s condition. This ensures alignment between the case outputs in the JF and the format of the training medical cases data. The systematic preprocessing of data has laid a solid foundation for subsequent accurate diagnosis and precise syndrome differentiation, significantly enhancing the model’s ability of syndrome differentiation-based treatment.
To enable JF with ability of precise syndrome differentiation, we particularly train and develop a TCM Syndrome Agent via fine tuning the LLM. Moreover, according to the final patient’s medical record from the multi-round consultation $Rec_{f}$, the TCM Syndrome Agent is used to identify the most likely syndrome type for the patient:
$TCM_{s}=Syn_{dif}(Rec_{f},\ Con_{t},\ prompt_{s},).$ (11)
For the detailed description of syndrome differentiation database preconditioning and the model fine tuning, see Sec.C.1 and Sec.C.2 in the appendix, respectively.
2.4 TCM Treatment Agent
In TCM diagnosis and treatment, treatment recommendations are essentially formulated from the perspective of syndrome differentiation. In addition, TCM experts also integrate various aspects of patient’s information, including chief complaints and medical history, dynamically adjusting treatment plans to ensure the precise and targeted treatment. This complicated and changeable procedure is a characteristic of TCM treatment but poses a challenge to the precise treatment of LLMs. To facilitate LLMs with accurate and targeted treatment recommendations ability, a Dual-Stage Retrieval Scheme (DSRS) is proposed. For more detailed information of this method, please refer to section D in the appendix.
According to DSRS, a TCM Treatment Agent that utilizes the patient’s medical record and syndrome differentiation result is developed to automatically screen and recommend the most suitable prescriptions from a structured TCM prescription database. It can be formulated as:
$TCM_{t}=Treat_{gen}(Rec_{f},DB_{TCM},prompt_{p}),$ (12)
where $DB_{TCM}$ represents the delicately constructed TCM prescription database containing various TCM fields, including internal medicine, gynecology, pediatrics, and surgery. Each entry in the database includes essential information such as disease categories, syndrome types, clinical manifestations, representative formulas, commonly used herbs, and other therapeutic methods. More information can be found in section F.4 in the appendix.
3 Experiments
In this section, we demonstrate the performance of JF on two core aspects: the accuracy of TCM syndrome differentiation and the comprehensiveness and effectiveness of clinical consultation, via a series of systematic experiments.
3.1 Datasets and Baseline Model
The dataset comes from three parts: the open-source TCM-SD [22] syndrome differentiation dataset, 48 modern TCM textbooks covering internal medicine, surgery, gynecology, and pediatrics, 700 ancient TCM texts. We manually cleaned the data and engaged three PhD students from different specialties at Beijing University of Chinese Medicine to proofread the data. We have constructed a high-quality TCM knowledge database for accurate treatment recommendation which consists of 80,000 prescription entries, including 1,000 classic formulas. Each entry contains key attributes such as disease type, syndrome type, clinical manifestations, syndrome mechanism, treatment methods, representative formulas, and commonly used medicines. We also have processed over 63,000 records of syndrome differentiation in raw data and selected 43,000 high-quality samples with detailed case descriptions and corresponding diagnoses for model training on syndrome differentiation.
To comprehensively evaluate the performance of JF, we selected several representative open-source TCM models, i.e., Zhongjing-7B [31], HuaTuo-32B [5], ShenNong-7B [28], and SunSimiao-7B [29], as baselines. Moreover, we incorporate four state-of-the-art (SOTA) LLMs i.e., Glaude-3.7 [2], DeepSeek v3 [10], Gemini-1.5-pro[23] and GPT-4o [18] for comparison. Various evaluation metrics are designed to ensure a comprehensive and reliable assessment. Different tasks adopts task-specific evaluation criteria, enabling a multidimensional analysis for the performance of JF.
3.2 Assessment of Syndrome Differentiation Accuracy
Based on the curated TCM syndrome differentiation data, the Qwen2.5-7B-Instruct model and the Roberta model are taken respectively as the foundation model of JF for training its syndrome differentiation ability. Here, syndrome differentiation can also be considered as a classification task, thus we train it via these two different types of models. The training was conducted on four A6000 GPUs, each with 48GB of VRAM. For the evaluation of syndrome differentiation accuracy, we randomly select 8,699 TCM cases as the test dataset, covering 170 different syndrome types. As the imbalance in sample sizes across different categories, we use weighted evaluation metrics to assess the model’s performance for the sake of fair and accurate evaluation. The metrics include weighted precision: $P_{w}$, weighted recall: $R_{w}$, and weighted F1: $F1_{w}$.
  Model $P_{w}$ $R_{w}$ $F1_{w}$     Model $P_{w}$ $R_{w}$ $F1_{w}$
JF-RoBERTa 0.8185 0.8230 0.8186     Glaude-3.7 0.2731 0.1206 0.1453
JF-Qwen-2.5-7B 0.8015 0.8145 0.8032     HuaTuo 0.3654 0.1735 0.1515
DeepSeek v3 0.4700 0.3032 0.3221     SunSimiao 0.3113 0.1497 0.1415
Gemini-1.5-pro 0.4532 0.1992 0.2209     ShenNong 0.0713 0.0945 0.0401
GPT-4o 0.4328 0.2138 0.2474     ZhongJing 0.0715 0.0598 0.0308
Table 1: Comparison of TCM syndrome differentiation performance across models.
As shown in Table 1, JF surpasses both the baseline models and the SOTA general models across all evaluation metrics in the TCM syndrome differentiation task. Specifically, JF model improves at least 74.1% in accuracy compared to the SOTA general model (DeepSeek-V3), and improves at least 124% in accuracy compared to the TCM model (HuaTuo-32B). Notably, the general models Glaude-3.7, DeepSeek v3, Gemini-1.5-pro and GPT-4o, due to their significantly larger number of parameters compared to the baseline models, outperformed other baseline models in certain metrics. However, their performance still does not reach the level of JF. These results demonstrate that our models possess the expert-level capability of syndrome differentiation. Here, we need to emphasize that the choice of foundation models is not a factor affecting its capabilities in syndrome differentiation, while the training method and data do. For more information, see the section 3.4.2
3.3 Medical Consultation Evaluation
For the evaluation of clinical consultation ability, we randomly selected 50 TCM medical cases from the database. These cases involve a variety of clinical conditions, which comprehensively reflect the diversity and complexity of TCM consultation. An LLM is used to act as the patient based on these cases, interacting with JF, baseline models and SOTA general models to accomplish the multi-round consultation and collect the patient’s medical information.
(a) Evaluation results for Proactivity
(b) Evaluation results for Accuracy
(c) Evaluation results for Practicality
(d) Evaluation results for Overall Effectiveness
Figure 3: Comparison of multi-round consultation capabilities across different models.
To comprehensively evaluate the multi-round medical consultation capabilities, 10 TCM experts were invited to assess the performance of each model across four key dimensions: Proactiveness, Accuracy, Practicality, and Overall Effectiveness. Pairwise comparisons are conducted between the proposed JF, each baseline and SOTA general model. Based on performance, each comparison was categorized as a win, loss, or tie (For detailed evaluation criteria of each dimension, refering to section B). The assessment in each dimension focused on the model’s performance in critical consultation skills, comprehensiveness and pertinence.
As shown in Figure 3, the MACCTM significantly improves the model’s overall performance on clinical consultation. Compared with various baseline and general models, the JF consistently outperforms all comparison models across the four dimensions of proactivity, accuracy, practicality, and overall effectiveness. These results indicate the MACCTM, which is a CoT-driven multi-agent collaborative mechanism, exhibits superior interaction and information processing capabilities during multi-round consultations. Based on performance, MACCTM provides a solid foundation for model to carry out accurate clinical diagnosis and highlights its potentiality in real-world applications.
3.4 Ablation Experiment
To assess the effectiveness of key components within the JF framework, we conducted an ablation study focusing on its multi-round consultation capabilities and dataset applicability. By systematically comparing different framework configurations and evaluating performance on real-world TCM cases, this experiment provides insights into the contributions of the TCM General Agent and the robustness of the constructed TCM syndrome differentiation dataset.
Figure 4: Performance comparison between JingFang with TCM General Agent and JingFang without TCM General Agent
3.4.1 Evaluation of TCM General Agent MACCTM
To evaluate the impact of the combination of TCM Specialist Agent and TCM General Agent in MACCTM on multi-round medical consultation and to analyze the key role and contributions of the TCM General Agent, this study compares two configurations: JF without TCM General Agent and JF with TCM General Agent, systematically analyzing their differences in performance in multi-round consultation. This experiment selects 100 representative TCM cases from medical data, ensuring typicality and diversity. To simulate multi-round interactions, we employ an LLM-based patient agent, which dynamically generates consultation responses from a patient’s perspective, effectively mimicking real-world consultation scenarios. To evaluate performance, we conducted a comparative analysis, assessing consultation results against the original case information. The evaluation focuses on the comprehensiveness (coverage of multi-dimensional health information) and the pertinence (adaptation to case-specific features). To ensure objectivity, GPT-4o is used as the evaluation tool to automatically compare consultation outputs from both configurations and score them based on these criteria.
As shown in Figure 4, the JF framework with TCM General Agent significantly outperforms the one without it in multi-round consultations. The evaluation favored the framework with the TCM General Agent 89 times for its comprehensiveness and pertinence, compared to only 11 times for the one without it, highlighting the critical role of the TCM General Agent in enhancing consultation quality. Additionally, the framework with the TCM General Agent achieves an average of 9.09 consultation rounds per patient, nearly doubling the 4.94 rounds of the one without it. This demonstrates that incorporating the TCM General Agent in the JF framework improves the completeness and coverage of multi-round consultations, and significantly reinforces the effectiveness and performance of the JF in medical consultations.
3.4.2 Applicability Assessment of the TCM Syndrome Differentiation Dataset
This ablation experiment is designed to demonstrate the effectiveness and applicability of our proposed method and constructed data for TCM syndrome differentiation and explore the influence of different foundation models (or the model with different size of parameters) on syndrome differentiation. We chose several representative open-source LLMs to test, which include the Qwen-2.5 series (Qwen-2.5-3B-Instruct, Qwen-2.5-7B-Instruct, Qwen-2.5-14B-Instruct) [30], the DeepSeek series (DeepSeek-7B-Chat) [9], and the Llama series (Llama3.1-8B-Instruct) [12], and the GLM series (GLM-4-9B Chat) [13], and the Gemma series (Gemma-3-12b-it) [24].
As shown in Figure 5, all models in the experiment exhibit advanced performance in TCM syndrome differentiation compared to their original version before being trained via our method and data. According to the results, we observe that the choice of foundation model is not the key factor for improving the syndrome differentiation ability, while the method and data do. Moreover, compared to the performance of GPT-4o (0.4328 in precison), the results highlight the potential applicability and value of the proposed TCM syndrome differentiation method and dataset for future research. For a detailed description of the ablation experiment, refer to Sec.E.
Figure 5: Effect of the proposed TCM syndrome differentiation method and dataset on different models
4 Conclusion and Discussion
[…]
Appendix
Appendix A Framework of JingFang
In this section, we introduce the framework of the proposed JingFang (referred to as the JF framework). As shown in Figure 1, following the authentic process of TCM clinical consultation, diagnosis, and treatment, JF framework is designed to consist of three main modules: TCM Consultation, TCM Syndrome Differentiation, and TCM Treatment Reconmmendation. It follows the systematically designed MACCTM and DSRS to fulfill the professional and accurate capabilities of consultation and syndrome differentiation-based treatment.
TCM Consultation Module mainly involves the following steps:
① Expert Team Construction: The Medical Record Agent summarizes the patient’s symptoms and basic information to generate the chief complaint. Based on the chief complaint, the relevant TCM Specialist Agents and a TCM General Agent are gathered to form the TCM expert team for the diagnosis.
② Consultation Construction: Each agent in the expert team constructs a consultation CoT based on their professional knowledge and the patient’s current medical condition, incorporating multiple follow-up questions to ensure the consultation process is logical and progressively in-depth.
③ Consultation Integration and Evaluation: Based on the consultation CoT provided by each agent, the Summary Agent integrates and evaluates the questions and key points within to form a summary of the consultation CoT.
④ Consultation Analysis and Optimization: Each agent in the expert team further analyzes the summary and proposes modification suggestions to the summarized consultation CoT, then sends them back to the Summary Agent until all the expert agents agree to officially begin the consultation.
⑤ Multi-Round Consultation: Based on the optimized consultation CoT, the TCM Consultation Agent conducts multiple rounds of consultation to further clarify the patient’s condition and collect vital information for diagnosis and treatment.
TCM Syndrome Differentiation Module is mainly for exact syndrome differentiation. Once sufficient key information is gathered from the consultation module, the trained TCM Syndrome Agent integrates and analyzes the data to determine the patient’s syndrome type for the accurate treatment of diseases.
TCM Treatment Recommendation Module provides professional and targeted treatment recommendations for patients. The Treatment Agent is generated through a DSRS to provide personalized treatment suggestions based on the patient’s syndrome type and the patient’s condition.
Appendix B TCM Consultation
To enable efficient consultation with the patient, we propose an "Intelligent Dynamic Consultation Information Collection and Termination" Algorithm. This algorithm automatically collects key points from the consultation CoT through a limited number of consultation rounds and automatically terminates the dialogue based on this consultation CoT when specific conditions are met. Based on TCM consultation knowledge, the key points are categorized into four types: symptom information, medical history information, lifestyle information, and other information, which helps to systematically extract crucial details. The algorithm’s inputs include the maximum consultation rounds, patient inputs, and the consultation CoT, and through these inputs, it generates a set of key points and drives the conversation forward. Finally, the algorithm outputs the collected set of medical information points.
The core logic of the algorithm ensures the consultation continues through an infinite loop until termination conditions are met. To ensure the effective collection of diagnostic information, the algorithm maintains the set of medical information points and updates this set according to the progress of the consultation, determining whether to terminate the dialogue. In addition, the algorithm defines two additional algorithms: one to obtain the set of key points to collect and the other to update the set of information points based on the patient’s input. These two auxiliary algorithms are executed through agents defined by prompts. The detailed methodology is presented in Algorithm 2.
Algorithm 2 Intelligent Dynamic Consultation Information Collection and Termination
1:  Initialization: $max\_rounds\leftarrow N$, $current\_round\leftarrow 0$, $medical\_points\leftarrow\emptyset$
2:  $cot\_points\leftarrow generate\_initial\_points(consultation\_cot)$
3:  while True do
4:     if $current\_round\geq max\_rounds$ then
5:        Terminate the loop.
6:     end if
7:     $patient\_input\leftarrow get\_user\_input()$
8:     if user opts to end the dialogue then
9:        Terminate the loop.
10:     end if
11:     $medical\_points\leftarrow update(medical\_points,patient\_input)$
12:     if $sufficient(medical\_points,cot\_points)$ then
13:        Terminate the loop.
14:     end if
15:     $current\_round\leftarrow current\_round+1$
16:  end while
Here, we detailed list the professional explanation of the four dimensions used in the evaluation of LLM’s consultation ability.
  1. 1.
    Proactiveness: Actively guiding and flexibly adjusting the direction of inquiry based on the initial symptom information provided by the patient, avoiding repetitive or ineffective questions, and ensuring the acquisition of accurate information.
  2. 2.
    Accuracy: The ability to pose questions that are closely related to the symptoms and precise, and to help patients express themselves accurately using clear language, thereby reducing misunderstandings and biases.
  3. 3.
    Practicality: Capable of comprehensively collecting the patient’s basic information (such as symptoms, lifestyle habits, etc.) and flexibly responding to different contexts, providing practical and feasible suggestions for clinical diagnosis and treatment.
  4. 4.
    Overall Effectiveness: Effectively integrating patient information, avoiding irrelevant distractions, providing valuable support for the diagnostic and treatment process, and enhancing diagnostic efficiency and effectiveness.
Appendix C TCM Syndrome Differentiation
C.1 Dataset Construction for Syndrome Differentiation
Figure 6: An example of the proposed universal TCM syndrome differentiation data cleaning method
During the construction of the syndrome differentiation dataset, we observed that TCM cases in the dataset often contain a substantial amount of information irrelevant to the patient’s condition, which cannot be easily removed through regularization techniques. For instance, phrases such as "The patient visited our hospital for further integrated traditional Chinese and Western medicine treatment" or "The patient was advised to undergo cranial CT and blood tests but refused" are not only extraneous to the core condition but also hinder the model’s ability to focus on key information about the illness. Moreover, the raw TCM case and diagnostic data frequently exhibit a high degree of redundancy in describing the patient’s condition and physical signs, resulting in a low density of valid information, which adversely impacts fine-tuning performance and the generalization ability of LLMs.
To address these issues, we proposed a universal syndrome differentiation data cleaning method based on the TCM framework of the "Ten Questions Song." This strategy aims to simulate the professional judgment of TCM experts using LLMs, accurately screening and reconstructing raw TCM case and diagnostic data to construct a high-quality syndrome differentiation dataset. Specifically, this approach strictly retains core information closely related to the patient’s condition, optimizes the medical value of the data, and organizes the content systematically according to the "Ten Questions Song" framework to ensure the dataset includes key elements required for TCM syndrome differentiation. Through this strategy, we effectively eliminate redundant information, increase the structuralization of the data, and better adapt it to the application requirements of LLMs in TCM syndrome differentiation tasks.
As shown in Figure 6, the data-cleaning process enforces strict adherence to the "Ten Questions Song" framework by applying explicit data extraction rules, retaining only information directly related to the patient’s condition. A comparison of the raw TCM case and diagnostic data with the cleaned dataset (Preconditioned TCM Data) reveals that the cleaned records are more focused on core symptoms, physical signs, and the evolution of the condition. Unrelated information is systematically removed, resulting in texts that are both precise and medically valuable. The core optimization points of this method include the following:
  1. 1.
    Removal of non-essential medical information: Exclude content unrelated to the diagnostic process, such as procedural details, examination recommendations, and results from modern medical instruments. Retain only symptoms and signs directly relevant to syndrome differentiation analysis.
  2. 2.
    Reduction of redundant descriptions and enhancement of information density: By simplifying repetitive records, the medical texts become more concise, improving data validity and utilization efficiency.
  3. 3.
    Strengthening the comprehensiveness and pertinence of syndrome differentiation dataset: Comprehensiveness is reflected in the inclusion of all key consultation dimensions from the TCM "Ten Questions Song," such as cold/heat, sweating, head/body, bowel movements, diet, sleep, medical history, and family history, ensuring the completeness of diagnostic criteria. Pertinence is achieved by prioritizing the retention of information related to core elements of syndrome differentiation (e.g., chief complaints, symptom changes, disease progression, and physical signs) while removing background or secondary information. This makes the dataset more aligned with the logic of TCM syndrome differentiation, thereby improving the accuracy of LLMs in clinical reasoning.
Overall, this data-cleaning strategy significantly enhances the quality of the syndrome differentiation dataset, providing reliable and efficient data support for TCM syndrome differentiation and large model training. It lays a solid foundation for the development of intelligent TCM diagnostic systems and serves as a reference paradigm for data optimization and model development in related research fields.
C.2 Model Training for Syndrome Differentiation
Figure 7: An example of TCM syndrome differentiation
Based on the syndrome differentiation data-cleaning method proposed in this study, a total of over 43,000 high-quality syndrome differentiation data entries were obtained. These entries were split into training and testing datasets at a ratio of 8:2. Subsequently, the Qwen2.5-7B-Instruct model was fine-tuned using the LoRA method. The form of the training dataset is shown in Figure 7. In the experiments, we employed four A6000 GPUs (each with 48GB of memory). After multiple rounds of experimental tuning, we determined the optimal training parameters to ensure the stable convergence of the loss function. The specific training parameters included: "lora rank": 64, "lora alpha": 16, "learning rate": "5e-4", and "batch size": 16.
Given that the syndrome differentiation task in this study can be treated as a classification problem, and Encoder-only architectures (such as BERT and RoBERTa) generally outperform Decoder-only architectures in classification tasks, we further trained lightweight syndrome differentiation models based on RoBERTa. In these experiments, a new classification layer was added to RoBERTa model, with its output dimension matching the number of task categories. The classification layer was trained from scratch, and fp16 mixed precision training was used to reduce memory usage and improve training efficiency. The training loss is shown in Figure 8.
Figure 8: Training loss of different models
Appendix D TCM Treatment Recommendation
Syndrome differentiation-based treatment is one of the core concepts in TCM. The fundamental idea is to identify the etiology and pathogenesis of a patient’s condition and then adopt a personalized treatment strategy. Different diseases may present similar syndrome differentiation; for instance, the syndrome of wind-cold attacking the lungs may manifest not only in the common cold but also in acute bronchitis. Therefore, experienced TCM practitioners often rely on the results of syndrome differentiation, integrating the patient’s medical history and constitution, to provide individualized treatment recommendations. This study employs a Dual-Stage approach to recall relevant TCM treatment knowledge based on the TCM formula dataset we have constructed. An example from the TCM formula database is shown in the Figure 9.
Figure 9: Example from the TCM Formula Database
In the first stage, the system screens candidate prescription data related to the patient’s syndrome type from the TCM prescription database based on the differentiation outcomes. This screening process utilizes the patient’s syndrome differentiation information, including syndrome type, etiology, and disease location, to narrow down the candidate prescriptions to those highly relevant to the patient’s condition, thereby ensuring the pertinence and effectiveness of subsequent treatment recommendations. In the second stage, the system employs an embedding method to vectorize the patient’s detailed medical history information and the clinical manifestations of the candidate prescriptions, subsequently calculating the similarity between them. This method allows for the assessment of how well the candidate prescriptions match the patient’s condition. Finally, based on similarity rankings, the top three cases most similar to the patient’s medical history information are selected (TOP-3). These "top three most similar cases" refer to cases that closely resemble the current patient’s condition across multiple dimensions, including syndrome differentiation information, medical history, and symptom presentation. The detailed methodology is presented in Algorithm 3. The "sparse" and "dense" embeddings in Algorithm 3 are two complementary vector representations in our Dual-Stage Retrieval Scheme: sparse embeddings (emb_sparse) are high-dimensional vectors with mostly zero elements, capturing exact lexical matches and keyword presence in medical records, excelling at precise identification of medical terminology and symptom descriptors; dense embeddings (emb_dense) are low-dimensional vectors with semantics distributed across all dimensions, capturing semantic relationships and contextual understanding beyond keyword matches, helping identify conceptually relevant prescriptions.
Algorithm 3 DSRS for TCM Treatment Recommendation
1:  Initialization:$max\_results\leftarrow 3$
2:  Preprocess Patient Medical Record: $cleaned\_record\leftarrow remove\_stopwords\_punctuation(medical\_record)$
[…]
5:  Execute Retrieval: $sparse\_results\leftarrow search(emb\_sparse,\ params)$ $dense\_results\leftarrow search(emb\_dense,\ params)$
6:  Hybrid Ranking with RRF Algorithm: $hybrid\_results\leftarrow rrf\_ranker(sparse\_results,\ dense\_results)$
7:  Output: $hybrid\_results$
We designed and conducted a comparative experiment to evaluate the performance differences between the DSRS and the direct symptom-based retrieval method. In this study, 100 clinical cases were randomly selected, and both the DSRS and single-stage direct symptom matching were applied under a unified recall threshold. The average similarity to the original data was then compared between the two approaches.
As shown in the Figure 10, the DSRS significantly outperformed the single-stage method in terms of similarity to the original cases. Further analysis revealed that the DSRS effectively reduced semantic noise introduced by non-target syndrome types that share similar symptom keywords, thereby improving the specificity and effectiveness of the retrieval. These findings demonstrate the potential of DSRS in TCM syndrome differentiation and treatment, contributing to more precise and intelligent recommendations in cases of “different treatments for the same disease” or “same treatment for different diseases.”
Figure 10: DSRS Ablation Experiment
Appendix E More Information of Ablation Experiment
E.1 Details of the Evaluation of TCM General Agent in Medical Consultation
To evaluate the impact of the combination of TCM Specialist Agent and TCM General Agent in MACCTM on multi-round medical consultation and to analyze the key role and contributions of the TCM General Agent, this study compares two configurations: JF without TCM General Agent and JF with TCM General Agent, systematically analyzing their differences in performance in multi-round consultation.
A total of 100 representative TCM cases were selected from real medical records to ensure data typicality and diversity, enhancing the clinical applicability of the experiment. The selection criteria included:
  1. 1.
    Information completeness: Each case must contain comprehensive medical history, including chief complaints, present illness, past medical history, and family history.
  2. 2.
    Diversity: The dataset includes various TCM syndrome types, ensuring that findings are applicable across different pathological conditions.
To maintain experimental control and standardization, we developed an LLM-based patient agent to simulate patient responses in multi-turn consultations. The patient agent was designed with the following constraints:
  1. 1.
    Strict adherence to case details to ensure responses align with actual patient conditions.
  2. 2.
    Concise and precise responses, avoiding unnecessary verbosity that could interfere with analysis.
  3. 3.
    Responses limited to doctor consultations, following real patient response patterns.
These constraints ensure stability and consistency in patient interactions, minimizing external variability. Figure 11 provides further details on the patient agent setup.
Figure 11: Patient Agent setup
To comprehensively assess the performance of different frameworks in multi-turn consultations, we used two core metrics:
  1. 1.
    Comprehensiveness (0-10 points): Evaluates whether the consultation covers key symptoms, medical history, and relevant factors, reflecting the physician’s understanding of the patient’s condition.
  2. 2.
    Pertinence (0-10 points): Assesses whether the consultation accurately focuses on the patient’s primary concerns and effectively guides subsequent diagnosis and decision-making.
To ensure objective scoring, we employed GPT-4o as an automated evaluation tool, conducting a quantitative analysis of consultation records generated by different frameworks and scoring them based on these metrics. Figure 12 presents the detailed scoring setup.
Figure 12: Evaluation Agent setup
To validate the significance of the experimental results, we performed an independent sample t-test and one-way ANOVA for statistical analysis. The results (Table 2) indicate that the JF framework with a TCM General Agent significantly outperforms the framework without a general agent in both comprehensiveness (8.91-9.0 vs. 6.164-6.406) and pertinence (7.981-8.099 vs. 5.222-5.478). The t-test and ANOVA results confirm statistical significance ($p<0.05$), demonstrating that the inclusion of a TCM General Agent effectively enhances the coverage and accuracy of multi-turn medical consultations, optimizing diagnostic strategies for physicians.
Label Value
p-value (t-test, Completeness) $1.34\times 10^{-113}$
p-value (t-test, Pertinence) $1.34\times 10^{-111}$
p-value (ANOVA, Completeness) $1.01\times 10^{-144}$
p-value (ANOVA, Pertinence) $3.12\times 10^{-133}$
Confidence Interval (JF, Completeness) (8.91, 9.0)
Confidence Interval (JF w/o, Completeness) (6.164, 6.406)
Confidence Interval (JF, Pertinence) (7.981, 8.099)
Confidence Interval (JF w/o, Pertinence) (5.222, 5.478)
Table 2: Statistical analysis of Completeness and Pertinence. JF represents JF with TCM General Agent, while JF w/o represents JF without TCM General Agent.
E.2 Applicability Evaluation of the TCM Syndrome Differentiation Dataset
To evaluate the applicability of the TCM syndrome differentiation dataset constructed in this study, we tested multiple open-source LLMs, including the Qwen-2.5 series (Qwen-2.5-3B-Instruct, Qwen-2.5-7B-Instruct, Qwen-2.5-14B-Instruct), the DeepSeek series (DeepSeek-7B-Chat), and the Llama series (Llama3.1-8B-Instruct).
Using LoRA, we fine-tuned these models with a training dataset comprising over 43,000 high-quality entries from the TCM syndrome differentiation dataset. The training was conducted on four A6000 GPUs, each with 48GB of VRAM. Key training parameters were set as follows: ‘"lora rank": 64‘, ‘"lora alpha": 16‘, ‘"learning rate": 5e-4‘, and ‘"batch size": 16‘.
Figure 13 illustrates the training loss curves for different models. The results show that all models exhibit a rapid decline in training loss during the initial phase, followed by a gradual stabilization as training progresses, indicating effective learning convergence. Among them, the Qwen-2.5-7B-Instruct and Qwen-2.5-14B-Instruct models achieve the lowest final loss values, suggesting superior adaptability to the dataset. The DeepSeek-7B-Chat and Llama3.1-8B-Instruct models show slightly higher final loss values, which may reflect their differing pretraining distributions and potential limitations in adapting to TCM-specific knowledge. Overall, these results confirm the dataset’s ability to facilitate efficient fine-tuning across various LLMs, with Qwen series models demonstrating the best performance in optimizing TCM syndrome differentiation.
Figure 13: Traing loss for different models
Appendix F Examples of Complete Medical Consultation
To clearly demonstrate the advantages of the JF framework proposed in this paper, four typical cases were selected, covering TCM surgery, TCM pediatrics, TCM gynecology, and TCM internal medicine. An in-depth analysis of the consultation processes, syndrome differentiation, and treatment recommendation was conducted.
F.1 Comprehensiveness of Multi-Round Consultations
The JF framework adopts a multi-round consultation approach to ensure that doctors can fully understand the patient’s chief complaints, medical history, symptom characteristics, and lifestyle habits. Through a structured consultation process, the framework allows for in-depth exploration of disease characteristics, avoiding the omission of critical information. For example:
  1. 1.
    In the TCM surgery case (Damp-Heat Sinking Downward Syndrome, as show in Figure 14), the consultation covered the degree of redness and swelling of ulcers, characteristics of exudate, pain nature, and disease progression to ensure accurate identification of the pathological mechanism.
  2. 2.
    In the TCM pediatrics case (Spleen and Kidney Yang Deficiency Syndrome, as show in Figure 15), the framework assessed multiple dimensions, such as diet, sleep, bowel and urinary habits, and immune status, for a comprehensive evaluation of the infant’s growth and development.
  3. 3.
    In the TCM gynecology case (Cold Congealing with Blood Stasis Syndrome, as show in Figure 16), the consultation not only focused on the characteristics of dysmenorrhea (e.g., pain location and alleviating factors) but also examined factors such as menstrual cycle, dietary habits, and bowel movements to clarify the etiology.
  4. 4.
    In the TCM internal medicine case (Yang Deficiency with Water Overflow Syndrome, as show in Figure 17), the consultation included the evolution of symptoms like palpitations, shortness of breath, and edema, combined with physiological indicators such as tongue coating and pulse condition, ensuring the accuracy of syndrome differentiation.
F.2 Pertinence of Multi-Round Consultation
The JF framework achieves pertinence information collection and personalized diagnosis and treatment planning through a progressive multi-round consultation approach. The pertinence of multi-round consultation in this framework is reflected as follows:
  1. 1.
    Structured Consultation Process: In the initial consultation phase, the doctor first collects the patient’s basic information and chief complaints. In the in-depth consultation phase, further inquiries about key symptom details are made based on patient feedback to ensure sufficient diagnostic evidence is obtained. For example, in the case of Cold Stagnation with Blood Stasis Syndrome, after hearing the patient’s description of pain, the doctor further inquires about the menstrual cycle, flow volume, and accompanying symptoms to ensure accurate syndrome differentiation.
  2. 2.
    Dynamically Adjusting the Consultation Focus: The doctor flexibly adjusts the focus of inquiry based on the patient’s responses. For instance, in the case of Damp-Heat Pouring Downward Syndrome, if the patient mentions the presence of induration around an ulcer, the doctor will further inquire about the ulcer’s formation process and any history of trauma to rule out other potential causes.
  3. 3.
    Comprehensive Syndrome Differentiation Using the Four Diagnostic Methods: The JF framework fully integrates inspection, listening and smelling, inquiry, and palpation in multi-round consultations to ensure a rigorous syndrome differentiation process. For example, in the case of Yang Deficiency with Water Overflow Syndrome, the doctor not only inquires about the location and severity of edema but also incorporates objective indicators such as tongue features (dim complexion, yellow-greasy coating) and pulse characteristics (deep and thin pulse) to further confirm the pathological mechanism.
F.3 Dynamic Optimization of consultation CoT
The JF framework incorporates dynamic optimization of CoT during consultations, allowing diagnoses to be continuously adjusted based on patient feedback. For example:
  1. 1.
    During the initial consultation phase, the framework collects basic patient information and asks about primary symptoms.
  2. 2.
    During the in-depth consultation phase, subsequent questions are adjusted based on patient responses to ensure sufficient diagnostic evidence is obtained. For instance, in the Cold Stagnation with Blood Stasis Syndrome case, after hearing the patient’s description of pain, the doctor further inquired about the menstrual cycle, flow volume, and accompanying symptoms to ensure accurate syndrome differentiation.
  3. 3.
    In the final diagnosis phase, the framework integrates multi-round consultation information, combines it with TCM theory for syndrome differentiation-based treatment and provides specific treatment plans.
F.4 Personalization of Treatment Recommendation
The JF framework not only focuses on disease diagnosis but also provides systematic and personalized treatment recommendations, encompassing herbal therapy, lifestyle adjustments, emotional regulation, and rehabilitation recommendations. For example:
  1. 1.
    In the TCM gynecology case, Ai Fu Nuan Gong Wan and Dang Gui Si Ni Tang were recommended to warm the meridians, dispel cold, activate blood, and relieve pain, with emphasis on keeping warm during menstruation and dietary adjustments.
  2. 2.
    In the TCM surgery case, Long Dan Xie Gan Tang was used to clear heat and detoxify, complemented by the external application of Huang Lian wash to promote ulcer healing.
  3. 3.
    In the TCM pediatrics case, Jian Pi Yi Shen Fang was employed to enhance the spleen and kidney functions of the infant, with additional recommendations for appropriate massage therapy to aid digestion.
  4. 4.
    In the TCM internal medicine case, Zhen Wu Tang was suggested to warm yang and promote water metabolism, alongside advice on keeping warm and engaging in moderate exercise to alleviate edema.
F.5 The Performance of JingFang in Authentic Clinical Scenarios
The application of the JF framework in multi-round TCM consultations fully demonstrates its advantages in comprehensiveness, pertinence, dynamic optimization, and personalized treatment plans. This framework not only enhances the accuracy of TCM diagnoses but also improves the syndrome differentiation-based treatment through the systematically designed and developed framework, showing significant value on practical application. Here are some examples of JF’s performance on medical consultation, syndrome differentiation and treatment in authentic clinical scenarios.
Appendix G Limitations
Despite JingFang’s expert-level performance in multi-round TCM consultation and syndrome differentiation, several limitations remain in real-world clinical settings. First, the system primarily relies on textual representations of tongue and pulse diagnostics, lacking the capability to directly process multimodal inputs such as tongue images and pulse waveform signals, thereby limiting its ability to fully emulate the “Four Diagnostic Methods” of TCM. Second, the training data contain substantial noise and irrelevant content, necessitating meticulous data cleaning and alignment to ensure model accuracy. Moreover, real-world consultations are highly complex, dynamic, and patient-specific. Although the current consultation COT framework based on the classical “Ten Questions Song” offers a degree of comprehensiveness and pertinence, it still falls short of fully capturing the diversity of real clinical scenarios. Future work will focus on incorporating multimodal perception mechanisms and developing models with stronger TCM diagnostic and reasoning capabilities to enhance generalizability and practical utility in authentic clinical environments.
Appendix H Broader Impacts
The development of JingFang marks a significant step toward the deep integration of TCM and LLMs, offering promising applications in primary care, remote consultations, and medical education. It also holds potential to facilitate the standardization and intelligent transformation of TCM diagnostic and treatment processes. In long-term health management tasks such as chronic disease management, JingFang can provide continuous and personalized TCM intervention recommendations, supporting patients in achieving more precise and effective recovery. However, the widespread deployment of such models also entails potential risks for instance, overreliance on AI systems may erode physicians’ clinical decision-making capabilities. Therefore, we advocate for a human-centered, assistive-first principle in practical deployment to ensure that the development of TCM-oriented LLMs is grounded in both patient safety and professional medical oversight.
Figure 14: An example of JF’s clinical performance in Damp-Heat Sinking Downward Syndrome differentiation and treatment
Figure 15: An example of JF’s clinical performance in Spleen and Kidney Deficiency Syndrome differentiation and treatment
Figure 16: An example of JF’s clinical performance in Cold Congealing and Blood Stasis Syndrome differentiation and treatment
Figure 17: An example of JF’s clinical performance in Yang Deficiency with Water Overflow Syndrome differentiation and treatment